OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

OpenStudy (anonymous):

Slope of a line passing through two points (x1,y1) and (x2,y2) = (y1-y2)/(x1-x2) Can you use this info ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Ok, solve for y.

OpenStudy (anonymous):

5-Y 7-6 = -3 5- y = 3

OpenStudy (anonymous):

so does y=8 or 2

OpenStudy (anonymous):

@shailkumar

OpenStudy (anonymous):

Actually, 5- y = -3 ==> y = 8